Consider a stock with, a price of 80, dividend growth 0.05, initial dividend 10 .

If the riskfree rate is 1% , what is the equity risk premium?

Enter as a rate with two decimal places, but do not type the % sign.

Please show your work!
Answer #1

Answer:

First of all we need to find cost of equity from the formula of dividend growth model of Gorden

Value of stock = D1 / (k - g)

D1 = Next year's expected dividend = Do(1+g)

Do = Current / Initial dividend

k = Cost of equity

g = dividend growth rate

80 = 10 (1.05) / ( k - 0.05 )

k - 0.05 = 0.13125‬

k = 0.13125‬ + 0.05

k = 0.18125‬ , 18.125 is equity cost

As per CAPM model

Ke = Rf + β (Rm - Rf)

Rf = Risk free return

β (Rm - Rf) = equity risk premium

18.125 = 1 + equity risk premium

equity risk premium = 18.125 - 1 = 17.125

Equity risk premium = 17.125